### v3

# Diffusing DeBias: Synthetic Bias Amplification for Model Debiasing

Abstract
Deep learning model effectiveness in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels. This results in a form of bias affecting training data, which typically leads to unrecoverable weak generalization in prediction. This paper aims at facing this problem by leveraging bias amplification with generated synthetic data: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing exploiting the inherent bias-learning tendency of diffusion models in data generation. Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which replace the original training set for learning an effective bias amplifier model that we subsequently incorporate into an end-to-end and a two-step unsupervised debiasing approach. By tackling the fundamental issue of bias-conflicting training samples memorization in learning auxiliary models, typical of this type of techniques, our proposed method beats current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
footnotetext: Correspondence: vito.paolo.pastore@unige.it
Figure 1: Comparison of bias amplifier training approaches. (a) Traditional methods that use real biased datasets lead to the memorization of biased-conflicting samples, resulting in suboptimal auxiliary models. (b) Our proposed approach leverages conditional diffusion models to learn class-specific biases and amplify them into synthetic images. Such generations can replace the original training set for learning an effective Bias Amplifier, eliminating memorization effects by avoiding exposure to real data.
1 Introduction
Deep learning models have shown impressive results in image classification, yet their success depends heavily on training data quality and representativeness. When a deep neural network is trained on a biased dataset, so-called “shortcuts”, corresponding to spurious correlations (i.e., biases) between irrelevant attributes and labels, are learned instead of semantic class attributes, affecting model generalization and impacting test performances. In other words, the model becomes biased towards specific training sub-populations presenting biases, known as bias-aligned, while samples not affected by the bias are referred to as bias-conflicting or unbiased.
In this scenario, several debiasing methods have been proposed [1, 2, 3, 4, 5], addressing the problem under different perspectives. Notably, the majority of available training samples are bias-aligned in a biased dataset (typically, $95\%$ or higher [6]). As such, common supervised debiasing strategies involve reweighting the training set to give more importance to the few available bias-conflicting data [5], for instance, by upsampling [7] or up-weighting [8] them. Conversely, unsupervised debiasing methods, which assume no prior availability of bias information, typically involve an auxiliary model to estimate and convey information on bias for reweighting the training set, employing their signal to guide the training of a debiased target model, while diminishing its tendency to learn bias shortcuts.
To this end, there are several existing works [6, 1, 5, 9] exploiting intentionally biased auxiliary models, trained to overfit bias-aligned samples, thus being very confident in both correctly predicting bias-aligned training samples as well as misclassifying bias-conflicting ones [5]. Such a distinct behavior between the two subpopulations is required to convey effective bias supervisory signals for target model debiasing, including gradients or per-sample loss values, which, e.g., should ideally be as high as possible for bias-conflicting and as low as possible for bias-aligned [6]. However, such desired behavior is severely hindered by bias-conflicting training samples, despite their low number. In fact, the low cardinality of this population impedes a suitable modeling of the correct data distribution, yet, they are sufficient to harm the estimation of a reliable bias-aligned data distribution, despite the latter having much higher cardinality. First, bias-conflicting samples in the training set act as noisy labels, affecting the bias learning for the auxiliary models [5]. Furthermore, these models tend to overfit and memorize bias-conflicting training samples within very few epochs, failing to provide different learning signals for the two subpopulations [10]. Despite available strategies, including the usage of the Generalized Cross Entropy (GCE) loss function [6, 3], ensembles of auxiliary models [1, 5], or even ad hoc solutions tailored to the specific dataset considered (e.g., training for only one epoch as in [11]), how to define a protocol for effective and robust auxiliary model training for unsupervised debiasing approaches remains an open issue.
Remarkably, recent works [3, 5, 1] underline how the target model generalization is profoundly impacted by the protocol used to train such an auxiliary model. In principle, if one could remove all bias-conflicting samples from the training set, it would be possible to train a bias-capturing model robust to bias-conflicting samples’ memorization and interference, and capable of conveying ideal learning signals [5].
To this aim, our intuition is to elude these drawbacks by skipping the use of the actual training data in favor of another set of samples synthetically generated by an adequately trained diffusion model. Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging fully biased synthetic data exploiting intrinsic properties of Conditional Diffusion Probabilistic Models (CDPMs). We aim to learn a per-class biased image distribution, from which we sample an amplified synthetic bias-aligned subpopulation. The resulting synthetic bias-aligned samples are then exploited for training a Bias Amplifier model instead of the original actual training set, solving the issue of bias-conflicting overfitting and interference by construction, as the bias-capturing model does not see the original training set at all. We refer to the resulting approach as Diffusing DeBias (DDB). DDB’s intuition is indirectly supported by recent works studying the problem of biased datasets for image generation tasks  [12, 13], and showing how diffusion models’ generations can be biased, as a consequence of bias present in the training sets (See Sec. 2).
This tendency can be seen as an undesired behavior of generative models, impacting the fairness of generated images, and recent papers have started to propose solutions for decreasing diffusion models’ inclination to learn biases [14, 13, 15]. However, we claim that what is considered a problem for generative tasks can be turned into a feature for image classification model debiasing, as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes, solving bias-conflicting training samples memorization, by construction (see Figure 1). For this reason, we refer to DDB as a plug-in for debiasing methods in image classification. To prove its effectiveness and versatility, we incorporate DDB’s bias amplifier into two different debiasing frameworks, named Recipe I and Recipe II, able to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO [8] algorithm (Recipe I, Sec. 3.3.1), and (ii) provide a loss function for the training set to be used within an end-to-end method (Recipe II, Sec. 3.3.2). Both approaches outperform the state-of-the-art on popular benchmark datasets. To summarize, our main contributions can be summarized as follows:
  • We introduce DDB, a novel and effective unsupervised debiasing framework that exploits intrinsic properties of CDPMs to learn per-class bias-aligned distributions in biased datasets, with unknown bias information. The resulting CDPM is used for generating synthetic bias-aligned images later employed for training a robust Bias Amplifier model, hence avoiding bias-conflicting training sample memorization and interference (Sec. 3).
  • We propose the usage of DDB as a versatile plug-in for debiasing approaches, designing two recipes: one based on a two-step procedure (Sec. 3.3.1), and the other as an end-to-end (Sec. 3.3.2) algorithm.
  • Both proposed approaches prove to be effective, surpassing state-of-the-art results by a significant margin (Sec. 4) across four different benchmark datasets.
2 Related Work
In this section, we provide a broad categorization of model debasing works.
Supervised Debiasing. Supervised methods exploit available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not). A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations [16, 17]. Alternatively, bias labels can be thought of as pre-defined encoded subgroups among target classes, allowing for the exploitation of robust training techniques such as G-DRO [8]. At the same time, bias information can be employed to apply regularization schemes aiming at forcing invariance over bias features [18, 19].
Unsupervised Debiasing. It considers bias information unavailable (the most likely scenario in real-world applications), thus having the highest potential impact among the described paradigms. Among such methods, a popular strategy consists in involving auxiliary models at any level of the debiasing process. We categorize these methods into two-step  [11, 9, 4]), and end-to-end  [6, 20]), in the following paragraphs.
End-to-end approaches. Here, bias mitigation is performed online with dedicated optimization objectives. Employing techniques such as ensembling and noise-robust losses have been proposed, though they require careful tuning of all the hyperparameters, which may require the availability of validation sets with bias annotations to work. In Learning from Failure (LfF), Nam et al. [6] assume that bias features are learned earlier than task-related ones: they employ a bias-capturing model to focus on easier (bias-aligned) samples through the GCE loss [21]. At the same time, a debiased network is jointly trained to assign larger weights to samples that the bias-capturing model struggles to discriminate. In DebiAN, Li et al. [20] jointly train an auxiliary model (the discoverer) and a classifier (the target model) in an alternate scheme so that the discoverer can identify bias learned by the classifier during training, optimizing an ad-hoc loss function for bias mitigation.
Two-step methods. Here, the auxiliary model is employed in a first step to identify bias in training data, generally through pseudo-labeling, while the second stage consists in actual bias mitigation exploiting the inferred pseudo-labels. In [11], the auxiliary model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples. Notably, they must rely on a validation set with bias annotations to stop the auxiliary model’s training after a few epochs to avoid memorization and to tune the most effective upsampling factor. In [1], a set of auxiliary classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample. Finally, a recent trend involves the usage of vision-language models to identify bias, further exploiting bias predictions to mitigate biases in classification models [22, 23, 24].
Among the described methods, a subset exploit the auxiliary model for bias amplification ([5, 6, 9]) under the general assumption that a bias amplification model would either misclassify or be less confident on bias-conflicting samples, with final debiasing performance strongly dependent on such behavior, and specifically, on how strongly the auxiliary model captures training bias [3, 5]. However, this fundamental assumption only holds if bias-conflicting samples are not memorized during training, which is likely to happen in very few training epochs [10, 3], especially considering that the very few training bias-conflicting samples are more likely to be overfitted, as reported in [8, 25]. One could argue that an annotated validation set could be used for properly regularizing the bias amplification model, but the latter cannot be assumed available a priori in real cases, this being en emerging argument of discussion among the community [10, 3].
For this reason, in this work and differently from the previously cited research, we propose that syntethic bias-aligned samples could be used instead of the original training set, solving the bias-conflicting memorization issue by construction. However, an unsupervised real-world scenario poses a problem for generating a pure distribution of bias-aligned samples. As a solution, we propose to leverage diffusion models for generating a synthetic bias-aligned training set to be used for training a bias amplifier, inspired by recent works, showing how they can manifest distorted, unfair, or stereotypical representations deviating from the true underlying data distribution [26], when exposed to biases present in the training data  [15]. While the vast majority of available works regarding generative models and debiasing refer to strategies for obtaining unbiased synthetic data generation [27, 28], up to our knowledge, we are the first to propose the usage of generative models for bias amplification. The closest available work, employing generative models for model debiasing, is Jung et al. [29], where the authors introduce a debiasing method exploiting a GAN [30] trained for style-transfer between several biased appearances, with a different objective. Here, a target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations.
In our work, we show that it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class, thus allowing the sampling of new synthetic images, sharing the same bias pattern of original train data. Our method provides a memorization-free bias amplification model, capable of providing precise supervisory signals for accurate model debiasing virtually compatible with any debiasing approach involving auxiliary models.
Figure 2: Schematic representation of our DDB framework. The debiasing process consists of two key steps: (A) Diffusing the Bias uses a conditional diffusion model with classifier-free guidance to generate synthetic images that preserve training dataset biases, and (B) employs a Bias Amplifier firstly trained on such synthetic data, and subsequently used during inference to extract supervisory bias signals from real images. These signals are used to guide the training process of a target debiased model by designing two debiasing recipes (i.e., 2-step and end-to-end methods).
3 The Approach
Our proposed debiasing approach is schematically depicted in Figure 2. In this section, we provide at first a general description of the problem setting (Sec.  3.1), and then, we illustrate in detail DDB’s two main components, which include bias diffusion (Sec. 3.2) and the two Recipes for model debiasing (Sec. 3.3).
3.1 Problem Setting
Let us consider a general data distribution $p_{\text{data}}$, typically encompassing multiple factors of variation and classes, and to build a dataset of images with the associated labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$ sampled from such a distribution. Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e.possible biases such as context, appearance, acquisition noise, viewpoint, etc.. In this case, data will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors. This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not presenting the same biases. In the same way, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$. Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting [13] In this context, biased and unbiased samples equivalently refer to bias-aligned and bias conflicting samples.. If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization. This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity [4, 8, 31]), which is dominating over the true target distribution semantics.
It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation [12]. For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in [12, 13]. While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature [6]), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
3.2 Diffusing the Bias
In the context of mitigating bias in classification models, the tendency of a CDPM to approximate the per-class biased distribution represents a key feature for training an auxiliary bias amplified model.
The Diffusion Process.
The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps [32]. Given a data point $\mathbf{x}_{0}$, the forward process adds Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, resulting in noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$. This forward process can be formulated for any timestep $t$ as:  $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ with $\alpha_{s}=1-\beta_{s}$. The reverse process then gradually denoises a sample, reparameterizing each step to predict the noise $\epsilon$ using a model $\bm{\epsilon}_{\theta}$:
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$ (1)
where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\sigma_{t}=\sqrt{\beta_{t}}$.
Classifier-Free Guidance for Biased Image Generation.
In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes and semantics. Classifier-Free Guidance (CFG) [33] introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
The CFG technique randomly omits conditioning during training (e.g., with probability $p_{\text{uncond}}=0.1$), enabling the model to learn both generation modalities. During the sampling process, a guidance scale $w$ modulates the influence of conditioning. When $w=0$, the model relies solely on the conditional model. As $w$ increases ($w\geq 1$), the conditioning effect is intensified, potentially resulting in more distinct features linked to $y$, thereby increasing fidelity to the class while possibly reducing diversity, whereas lower values help to preserve diversity by decreasing the influence of conditioning. The guided noise prediction is given by:
$\bm{\epsilon}_{t}=(1+w)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)-w\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t),$ (2)
where $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)$ is the noise prediction conditioned on class label $y$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction. This modified noise prediction replaces the standard $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ term in the reverse process formula (Equation 1). In this work, we empirically show how CDPM learns and amplifies the underlying biased distribution when trained on a biased dataset with strong spurious correlations, allowing bias-aligned image generation.
3.3 DDB: Bias Amplifier and Model Debiasing
As stated in Sec. 2, a typical unsupervised approach to model debiasing relies on an auxiliary intentionally-biased model, named here as Bias Amplifier (BA). This model can be exploited in either 2-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
3.3.1 Recipe I: 2-step debiasing
Figure 3: Overview of Recipe I’s 2-step debiasing approach.
The adopted 2-step approach consists in 1) applying the auxiliary model trained on biased generated data to perform a bias pseudo-labeling, hence estimating bias-aligned/bias-conflict split of original actual data, and 2) apply a bias supervised method to train a debiased target model for classification. For the latter, we use the group DRO algorithm [8] (G-DRO) as a proven technique for the pure debiasing step. In other words, being in the unsupervised bias scenario where the real bias labels are unknown, we estimate bias pseudo-labels performing an inference step by feeding the trained BA with the original actual training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified. Among possible strategies to assign bias pseudo-labels, such as feature-clustering [9] or anomaly detection [3], we adopt a simple heuristic based on the BA misclassifications. Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate bias-conflicting samples as
$\hat{c}_{i}=\mathds{1}\left(\hat{y}_{i}\neq y_{i}~{}\land~{}\mathcal{L}(\hat{y}_{i},y_{i})~{}>~{}\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})\right)$ (3)
where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$. Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Eq. 3). Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 3.
The above filtering operation refines the plain error set, restricting bias-conflicting sample selection to the hardest training samples, with potential benefits for the most difficult correlation settings ($\rho>0.99$). Later in the experimental section, we provide an ablation study comparing different filtering ($\gamma$) configurations and plain error set alternatives.
3.3.2 Recipe II: end-to-end debiasing
A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one [6] or more [1, 5] auxiliary intentionally-biased models. Here, we design an end-to-end debiasing procedure, denoted as Recipe II, incorporating our BA by customizing a widespread general scheme, introduced in the Learning from Failure (LfF) method [6]. LfF leverages an intentionally-biased model trained using Generalized CE (GCE) loss to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score. Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 4. Such loss function is used to obtain a weighting factor for the target model loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$. Coarsely speaking, $r$ should be low for bias-aligned and high for bias-conflicting samples.
Figure 4: Overview of Recipe II’s end-to-end debiasing approach.
4 Experiments
In this Section, after the description of the benchmark datasets employed in our work, we provide general implementation and training details, followed by a discussion of the obtained results and a thorough ablation analysis.
In the Supplementary Material, we report Recipe I and Recipe II full implementation details, more examples of generated images, and a quantitative analysis of the dataset bias captured from the CDPM with the assistance of a Vision-Language Model.
Figure 5: Examples of synthetic bias-aligned images across multiple datasets. Each grid shows synthetic images for specific classes across biased datasets, revealing how the model aligns with dataset-specific biases in different contexts.
4.1 Datasets
Our experiments utilize five distinct datasets: Waterbirds, Biased Flickr-Faces-HQ (BFFHQ), Biased Action Recognition (BAR), and ImageNet-9/ImageNet-A. Waterbirds  [8] is an image dataset exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs. land). It has 4,795 training images, 1,199 images for validation, and 5,794 for testing. BFFHQ [4] builds upon FFHQ [34] by introducing demographic biases for face recognition, while BAR  [6] is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts. For these two last datasets, we specifically refer to their original versions introduced respectively in [6] and [4]: BFFHQ has a total of 19,200 training images where  ( $y\in\{\text{young, old}\},~{}~{}b\in\{\text{female, male}\},~{}\rho=0.995$). Then, it provides 1,000 validation images (all bias-aligned) and 1,000 testing images with uniformly distributed labels and bias attributes. BAR presents 6 different classes, but does not provide explicit bias annotations or a validation set, with 1,941 images for training and 654 for testing. Finally, ImageNet-A consists of 7,500 real-world images from a 200-class subset of ImageNet, where deep learning models consistently struggle to make correct predictions. Originally introduced by Hendrycks et al. [35] to evaluate adversarial robustness, it is also employed as a model debiasing benchmark, usually as a bias-conflicting evaluation set for models trained on ImageNet-9 [36], a collection of images with 57,200 samples where categories are super-classes of ImageNet, put together to exhibit textural and shape biases [36]. As such, both our CDPM and Bias Amplifier are trained on ImageNet-9. Then, we evaluate both recipes on ImageNet-A in terms of Average Accuracy, following the same protocol of [36, 1, 31]
4.2 Implementation Details
In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels. For computational efficiency, input images are resized to $64\times 64$ pixels. Input images are normalized with respect to ImageNet’s mean and standard deviation. The CDPM is trained from scratch for $70,000$, $150,000$, $200,000$, $100.000$ iterations for the datasets BAR, Waterbirds, BFFHQ, and ImageNet-9 respectively. Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$). Optimization is performed using MSE loss and AdamW optimizer, with an initial learning rate of $1e^{-4}$, adjusted by a CosineAnnealingLR [37] scheduler with a warm-up period for the first $10\%$ of the total iterations. As per the Bias Amplifier training, we use synthetic images from CDPM. A Densenet-121 [38] with ImageNet pre-training is employed across all datasets, featuring a single-layer linear classification head. The training uses regular Cross-Entropy loss function and the AdamW optimizer [39]. For BFFHQ and BAR, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds, and ImageNet-9. Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs. We apply standard basic data augmentation strategies (following [4, 6]), including random crops and horizontal flips. For our debiasing recipes, we rely on the same backbones used in the existing literature, i.e.ResNet-18 for BFFHQ, BAR, ImageNet-9/A, and a ResNet-50 for Waterbirds [6, 4, 23]. In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments. Finally, similarly to [20, 6, 3], we do not rely on a validation set for regularization. This choice is related to real-world application suitability, where no guarantees exist regarding the degree of spurious correlations present in a held-out subset of data used as validation, possibly resulting in detrimental regularization [10].
4.3 Results
4.3.1 Synthetic bias-aligned image generation
First, we present qualitative image generation results across multiple datasets (see Fig. 5), to validate the effectiveness of our bias diffusion approach. The generated samples maintain good fidelity and capture the typical bias patterns of each training set per class, confirming our conditioned diffusion process’s ability to learn and diffuse inherent biases. This is confirmed by screening the generated images by 2 independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample. In addition, we also carried out an objective evaluation of the generated image content by adopting an image captioner, specifically BLIP-2 [40]. We estimate a caption for each image per class and count the frequency of every word. Suppl. Material (Sec. A.E) shows a histogram of the detected keywords related to the Waterbirds dataset, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced. The same behavior also emerges in other datasets used in this study, where a consistent description of the bias attributes can be noted, as reported in Sec. A.E. These results highlight our CDPM’s capability to generate synthetic bias-aligned data, setting the stage for our subsequent debiasing techniques.
4.3.2 Classification accuracy
Table 1 reports the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets. We report Worst Group Accuracy (WGA) for Waterbirds, as in [8, 5]. For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following [4, 20]. BAR and ImageNet-A do not provide bias-attribute annotations, hence, we report just the average test accuracy. For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively. It’s worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
Method Unsup Waterbirds BAR BFFHQ ImageNet-A
WGA Avg. Avg. Confl. Avg.
[…]
LISA [41] $89.20$
[…]
Park et al. [44] $71.68$
[…]
CDvG+LfF [29] $84.80$ $62.20\text{\scriptsize{$\pm$0.45}}$ $34.60$
[…]
DDB-I
(w/ plain err. set)
[…]
Table 1: Results of DDB Recipes I and II on benchmark datasets. Ticks () indicate that the method is bias-unsupervised. We highlight in bold the best results from unsupervised debiasing works, second-best are underlined.
Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results (see Table 1). Notably, in Waterbirds, DDB’s WGA is higher than G-DRO supervised (91.56% vs. 91.40% of G-DRO [8]) and all the other unsupervised methods’ ones. DDB-I is providing slightly worse performance with respect to DDB-II, but still better than the state of the art. In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN [20]), in terms of conflicting accuracy. In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to ETF-Debias [43] (73.60%), the best state-of-the-art results to date, and provides a boost of 7.45% over the recent DeNetDM [45].
4.4 Ablation analysis
In this section, we provide an extensive ablation analysis of DDB main components. Without losing generality, all ablations are performed on Waterbirds and exploit Recipe I for model debiasing. Additional ablations on Recipe II can be found in the Suppl. Material. The reported analyses regard several aspects of the data generation process with respect to the Bias Amplifier, the effects of the filtering option in Recipe I, and the behavior of DDB on unbiased data.
Bias amplifier data requirements. All our main results are obtained by training a BA on $1000$ synthetic images per class. Here, we want to measure how this choice impacts the effectiveness of our Bias Amplifier. From Table 2 (left), we observe that even 100 synthetic images are enough for good debiasing performance despite not being competitive with top accuracy scores. As we increase the number of training images, WGA improves, reaching the best performance for $2000$ images per class.
CFG strength and generation bias. The most important free parameter for the CDPM with CFG is the guidance conditioning strength $w$. Such strength can affect the variety and consistency of generated images. Here, we investigate DDB dependency on this parameter. Table 2 (right) shows a limited impact on debiasing performance. WGA variations along different $w$ values are limited, with the best performance obtained for $w=1$.
# Synth. images for BA training
# Imgs. $100$ $200$ $500$ $1000$ $2000$
WGA $86.25$ $87.10$ $87.67$ $89.05$ $\bf 90.81$
Guidance strength
$w$ 0 1 2 3 5
WGA 87.85 90.81 89.25 89.56 88.32
Table 2: Ablations on: (left) synthetic image count for training the BA on Waterbirds; (right) Impact of guidance strength $w$ on CDPM sampling, evaluated on Waterbirds, employing Recipe I.
Group extraction via BA plain error sets. In Sec. 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution. In Table 1, the last row (DDB-I w/ plain err. set) reports our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier. Our results show a trade-off between the two alternatives. The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings ($\rho=0.950$) (e.g., Waterbirds), gaining an advantage from such a coarser selection. On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like BFFHQ ($\rho=0.995$).
Bias-Identification Accuracy To frame the goodness of our Bias Amplifier in correctly identifying aligned and conflicting training samples, we provide the accuracy of such process when considering samples misclassified by the BA. Specifically, we report a bias-identification accuracy of $89.00\%$ for Waterbirds and $96.00\%$ on BFFHQ. Our very high accuracy supports the effectiveness of using the proposed protocol, which is not dependent from any careful regularization, bias-annotated validation sets, or large ensembles of auxiliary classifiers.
DDB impact on unbiased dataset. In real-world settings, it is generally unknown whether bias is present in datasets or not. Consequently, an effective approach, while mitigating bias dependency in the case of a biased dataset should also not degrade the performance when the bias is not present. This ablation study aims to assess DDB from this perspective. To obtain this insight, we apply the entire DDB pipeline (adopting Recipe I) on a common image dataset such as CIFAR-10  [46], comparing it with traditional ERM training with CE loss. In this controlled case, Recipe I provides a test accuracy of $84.16\%$, with a slight improvement of $\sim 1\%$ compared to traditional ERM ($83.26\%$) employing the same target model (ResNet18). This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard learning scenarios, potentially favoring the learning of the most challenging samples.
5 Conclusions
In this work, we present Diffusing DeBias (DDB), a novel debiasing framework capable of learning the per-class bias-aligned training data distribution, leveraging a CDPM. Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased target model. The usage of synthetic bias-aligned data permits avoidance of detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model. DDB is effective and versatile, acting as a plug-in for unsupervised debiasing methods. In this work, we design a two-step and an end-to-end debiasing recipes incorporating DDB’s Bias Amplifier: they outperform state-of-the-art works by a margin on four popular biased benchmark datasets, including the most recent relying on vision-language models [23, 47]. Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable for debiasing in real-world applications.
Limitations. DDB’s main limitations rely on the well-known high computational complexity of diffusion models. As a trade-off between computational efficiency and generation quality, we reduce the training sets resolution to $64\times 64$ when training the CDPM. Still, our CDPM requires $\sim 14$ hours for training on the largest considered dataset (i.e., BFFHQ, having $19,200$ images) on an NVIDIA A30 with $24$ GB of VRAM. Another limitation inherited from diffusion models is that the generation quality may decrease if the training dataset is too small, impacting DDB applicability for very small-scale datasets.
Appendix
Appendix A Ablation Studies with DDB’s Recipe II
This appendix section focuses on additional ablation studies, dedicated to our Recipe II. This set of ablations aims at providing additional evidence of Recipe II’s robustness and effectiveness, similar to what is shown in Sec. 4.4 of the main paper regarding Recipe I. As in the main paper, we run all our ablations on Waterbirds.
Bias amplifier data requirements.
# Synth. images for BA training
# Imgs. $100$ $200$ $500$ $1000$ $2000$
WGA $84.42$ $90.03$ $90.18$ $90.78$ $\bf 91.56$
Table 3: Ablation on the number of training images for the Bias Amplifier, in terms of WGA on Waterbirds, when using DDB’s Recipe II.
This ablation study measures how the number of training images for the Bias Amplifier impacts DDB’s Recipe II. Results are evaluated in terms of WGA on Waterbirds (see Table 2). $200$ images ($100$ images per class) are enough to obtain state-of-the-art competitive results (WGA equal to $90.03$), with an increasing number of images corresponding to higher WGA values, up to a maximum of $91.56$ for $2000$ total images.
CFG strength and generation bias.
Guidance Strength
$w$ 00 $\bf 1$ $2$ $3$ $5$
WGA $87.07$ $\bf 91.56$ $88.01$ $89.87$ $88.30$
Table 4: Impact of guidance strength $w$ on the final WGA on Waterbirds, when using DDB’s Recipe II.
Here, we evaluate DDB’s Recipe II dependency on the guidance strength parameter. Results are summarized in Table 4. The best results correspond to the usage of $w=1$, with different guidance strength parameter values having a relatively limited impact on debiasing performance.
DDB impact on an unbiased dataset.
Similarly to what is shown in Sec. 4.4, we want to ensure that applying Recipe II on an unbiased dataset does not degrade the final performance. Also Recipe II demonstrates to be applicable in unbiased scenarios, as it measures a final test accuracy of $83.87\%$ on CIFAR10, which is slightly higher than the ERM baseline ($83.26\%$).
Appendix B Full Implementation Details of Recipe I (two-step)
For the hyperparameters of Recipe I, we use for each dataset the following configurations. BAR: batch size = 32, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. Waterbirds: batch size = 128, learning rate = $5\times e^{-4}$, weight decay = 1.0, training for 60 epochs. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 1.0, training for 60 epochs. The GDRO optimization is always run with the --robust and --reweight_groups flags.
Appendix C Full Implementation Details of Recipe II (end-to-end)
For BAR, we use the same parameters reported in [6]: batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50. Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs. For these recipes, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
Appendix D Additional Biased Generations
In this Section, we present $100$ additional per class generated images for each dataset utilized in the study, as produced by our CDPM. In the Waterbirds dataset (Figure 6), the model adeptly captures and replicates the correlation between bird species and their background. In the BFFHQ dataset (Figure 9), it effectively reflects demographic biases by mirroring gender-appearance correlations. The results from the BAR dataset (Figure 8) exhibit the model’s adeptness in handling multiple bias patterns, preserving typical environmental contexts for different actions. Finally, the ImageNet-9 dataset (Figure 7) stereotypes the various classes and their backgrounds well. Overall, the model generates high-quality samples that are fidelity-biased across all datasets. Our research confirms that the model exhibits bias-learning behavior, as it not only learns but also amplifies existing biases within diverse datasets. We leverage this characteristic to produce synthetic biased data aimed at enhancing the robustness of debiasing techniques.
Figure 6: Synthetic bias-aligned image generations for each class of the Waterbirds dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
Figure 7: Synthetic bias-aligned image generations for each class of the BAR dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
Figure 8: Synthetic bias-aligned image generations for each class of the ImageNet-9 dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
Figure 9: Synthetic bias-aligned image generations for each class of the BFFHQ dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
Appendix E Bias Validation via Captions in Synthetic Images
As an ulterior confirmation of bias in CDPM-generated images, besides two independent annotators, we implement an unsupervised analysis pipeline using the pre-trained BLIP-2 FlanT5XL model [40] for zero-shot image captioning without any task-specific prompting, for avoiding any guidance in the descriptions of the synthetic data. Our method processes the resulting captions through stop word removal and frequency analysis to reveal underlying biases without relying on predetermined categories or supervised classification models. Figure 10 presents the keyword frequency distribution for the Waterbirds dataset, where the most prevalent terms naturally correspond to known bias attributes (e.g., environmental contexts). The same is true for the keyword frequency distribution of the BFFHQ dataset (Figure 11) and BAR (Figure 12) where the most prevalent terms are related to the gender and the context respectively. These findings confirm that generated images are bias-aligned. Furthermore, the obtained word frequency histograms suggest that images generated by DDB’s bias diffusion step may provide information useful to support bias identification (or discovery) at the dataset level.
Figure 10: Word frequency analysis from 1000 generated captions for Waterbirds classes ‘landbird’ (left) and ‘waterbird’ (right), showing top 10 most frequent terms.
Figure 11: Word frequency analysis from 1000 generated captions for BFFHQ classes ‘young’ (left) and ‘old’ (right), showing top 10 most frequent terms.
(a) Class 0
(b) Class 1
(c) Class 2
(d) Class 3
(e) Class 4
(f) Class 5
Figure 12: Word frequency analysis from 1000 generated captions for BAR classes, showing top 10 most frequent terms for each class.